OpenStudy (anonymous):

Please help! Solve (3x^2+13x-10) ÷ (x+5)

OpenStudy (campbell_st):

well I'd factor the quadratic equation... then look for a common factor.

OpenStudy (anonymous):

@campbell_st i need help solving the whole thing because i dont know how to solve it.

OpenStudy (campbell_st):

so you can't factor the quadratic.... forget solving at this stage... the 1st task is factoring...

OpenStudy (anonymous):

is it 3x^25 divided by x+5

OpenStudy (campbell_st):

here is a lesson if factoring difficult quadratics in the form \[ax^2 + bx + c\] multiply a and c... so in your question its 3 * -10 = -30 now find the factors of -30 that add to 13.... the larger factor is positive and the smaller is negative... any thoughts.

OpenStudy (anonymous):

im sorry i have no idea. i never did these problems bofore

OpenStudy (campbell_st):

ok.... so the factors of -30 are 15 and -2 next write the factored form as (ax + factor 1)(ax + factor2) ------------------------ a so in your question its \[\frac{(3x + 15)(3x -2)}{3}\] take out common factors in the numerator \[\frac{3(x + 5)(3x - 2)}{3}\] cancel the common factor and its \[3x^2 + 13x -10 = (x +5)(3x -2)\] so now the initial problem becomes \[(x + 5)(3x -2) \div(x + 5)~~~or~~~~\frac{(x + 5)(3x -2)}{(x + 5)}\] now you can get a solution

OpenStudy (anonymous):

how do i get the solution?